OpenStudy (anonymous):

What is the exact distance from (–4, –2) to (4, 6)? (4 points) units units units units

OpenStudy (mvanessa):

HERE ARE THE ANSWER CHOICES \[\sqrt{24} units\] \[\sqrt{88} units\] \[\sqrt{100} units\] \[\sqrt{128} units\]

OpenStudy (mvanessa):

AND THE ANSWER IS THE FOURTH ONE

OpenStudy (mvanessa):

WHICH IS